OpenStudy (anonymous):

GIVE MEDAL Edward plans to put $75 into a savings account. He can place his money into an account represented by b(x) = 7x + 75, or into another account represented by c(x) = 75(1.06)x. Which account has the highest value in 2 years? Which account has the highest in 12 years? c(x) has the highest value in 2 years; c(x) has the highest value in 12 years b(x) has the highest value in 2 years; b(x) has the highest value in 12 years c(x) has the highest value in 2 years; b(x) has the highest value in 12 years b(x) has the highest value in 2 years; c(x) has the highest value in 12 years

OpenStudy (anonymous):

and it is not option C

OpenStudy (anonymous):

Let me work it out now

OpenStudy (error1603):

Still here?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I think c(x)=75(1.06)^x is greater

OpenStudy (anonymous):

Im going to show you a screenshot because their is exponents

OpenStudy (anonymous):

I think it is because I got a higher number when I plugged in 2 and 12 c(x) has the highest value in 2 years; c(x) has the highest value in 12 years

OpenStudy (error1603):

So the correct answer would be C I got C

OpenStudy (anonymous):

Its not C I chose that one and got it incorrect

OpenStudy (anonymous):

this is the 3rd time im doing this question

OpenStudy (error1603):

Wait let me try again

OpenStudy (texaschic101):

b(x) = 7x + 75 c(x) = 75(1.06)^x b(2) = 7(2) + 75 c(2) = 75 (1.06^2) b(2) = 14 + 75 c(2) = 75 * 1.1236 b(2) = 89 c(2) = 84.27 b(12) = 7(12) + 75 c(12) = 75(1.06)^12 b(12) = 84 + 75 c(12) = 75 * 2.012 b(12) = 159 c(12) = 150.9

OpenStudy (error1603):

In two years b(x)=89 C(x)=159 In 12 years b(x)=159 C(x)=954

OpenStudy (anonymous):

So it would be b(x) has the highest value in 2 years; b(x) has the highest value in 12 years?

OpenStudy (anonymous):

c(x) has the highest value in 2 years; c(x) has the highest value in 12 years b(x) has the highest value in 2 years; b(x) has the highest value in 12 years

OpenStudy (anonymous):

Is it the first or second?^^^^^^^^^

OpenStudy (error1603):

That's as close as we can get try b(x)

OpenStudy (anonymous):

Ok ima try the second option out of them two

OpenStudy (error1603):

Ok

OpenStudy (texaschic101):

bx has highest value in 2 yrs and in 12 yrs

OpenStudy (anonymous):

I have another question.. can yall help me figure it out?

OpenStudy (error1603):

Sure

OpenStudy (texaschic101):

is it the one you posted a screenshot for ?

OpenStudy (anonymous):

The one I posted a screenshot for you answered @texaschic101

OpenStudy (anonymous):

This is the other difficult one im confused about.

OpenStudy (texaschic101):

the one you posted a screenshot for is the same one that you just posted again...lol 2^6 > 4(6) + 12 64 > 24 + 12 64 > 36 2^5 > 4(5) + 12 32 > 20 + 12 32 > 32....incorrect so it is 6 months....anything less will not work

OpenStudy (error1603):

umm

OpenStudy (anonymous):

Yes thats correct @texaschic101 I was just working it out and thats what I got :)

OpenStudy (anonymous):

Thank You Ill give you a medal once again!

OpenStudy (error1603):

k